OpenStudy (anonymous):

Find the equation of the tangent line to the given curve at the point (0,9). y=4x^2+8x+9

OpenStudy (saifoo.khan):

First step, find the dy/dx @ x = 0 4x^2 + 8x + 9 dy/dx = 8x + 8 @ x= 0 8(0) + 8 8 Therefore, slope = 8

OpenStudy (anonymous):

first we are going to find the derivative of our function which is y' = 8x+ 8 then we are going to let x=0 in our slope equation y' so y'=8 and that is our slope and now we have everything we need to find the tangent line m=8 x=0 y=9 let us use this equation where (y -y1) =m(x-x1) and therefore y-9=8(x-0) which is y=8x + 9:)

OpenStudy (saifoo.khan):

Next, Use the point given, slope and (y -y1) =m(x-x1) to work out.

OpenStudy (anonymous):

thank you :) life savers lol